Transverse waves travel through a string where the tension equals 13.89 N with a speed of 24.54 m/s. What tension would be required for a wave speed of 28.34 m/s?


Round your answer to 2 decimal places


1
Expert's answer
2021-04-26T17:03:54-0400

Express the velocity on terms of tension and mass per unit length:


"v=\sqrt{\frac T\mu},\\\space\\
\mu=\frac{ T}{v^2}."

Since we use the same spring, the mass per unit length is the same, which means that the following tension is required to get another velocity "u":


"T_a=\mu u^2=T\frac{u^2}{v^2}=18.52\text{ N}."